# revtex4-2Repair the float

### v1

@@ VI.3 Operation Speed
First of all, the presented method refers to the equilibrium type of quantum thermometry, meaning that the population measurement is performed after the qubit thermalization to the bath, which happens on time scales not shorter than the relaxation time $\tau_{1}$. In the presented experiment the time delay between the consecutive measurements, which is necessary for the qubit relaxation after the readout and thermalization of the qubit, was $\sim 10\,\tau_{1}$, which takes up most of the measurement time, and defines the speed of the response of the system. Pulse application and the qubit state readout can be limited to sub-microsecond range. Finally, the total measurement time is defined by the averaging number $t_{total}\sim N_{av}\cdot 10\,\tau_{1}=10^{6}\tau_{1}$, with $N_{av}\simeq 10^{5}$, meaning a time scale of seconds or minutes per point, depending on $\tau_{1}$.
VII Conclusion
@@
In this work we experimentally realized equilibrium quantum thermometry of the cryostat temperature, based on the population distribution measurements of the three lowest energy levels of a transmon qubit. The presented technique allowed for temperature measurements in the range of $60-220\,$mK, in which the qubit population follows the Maxwell-Boltzmann distribution. Analysis of temperature dependencies of the qubit relaxation rate and the photon number showed that the qubit was thermalized to the cryostat. The working temperature range was defined by saturation of the effective qubit temperature in the low temperature limit and the quasiparticle suppression of the qubit relaxation time above $200$ mK. We compared the experimental data to a numerical model, describing time evolution of the qubit population, showing good agreement. We believe, a similar approach could be exploited for quantum thermometry of various on-chip structures and for studies of coupling between quantum objects to different thermal baths. Finally, this experimental platform allows for exploitation of quantum thermometry algorithms including nonequilibrium techniques for faster measurements and dynamical control of the thermometers for enhanced precision.
@@ Acknowledgements.
The authors are grateful to Gershon Kurizki for fruitful discussions and Yu-Cheng Chang for assistance with measurements and fabrication. We acknowledge the provision of facilities and technical support by Aalto University at OtaNano - Micronova Nanofabrication Centre and OtaNano - Low Temperature Laboratory.
@@ Appendix A Relaxation and Excitation Rates for Coupling to a Thermal Bath
where $n=1/\left[1-\exp\left(-\hbar\omega_{0}\beta\right)\right]$ is the Bose-Einstein distribution. The total decay rate in this case would be
@@ A.1 Effective Temperature of a Two-Level System
Let us consider a qubit with the ground state |$\ket{g}$ and excited state |$\ket{e}$, thermalized with bath with temperature $T$. Then the population ratio of the qubit levels should follow the detailed balance principle $p_{e}/p_{g}=e^{-\hbar\omega\beta}=\Gamma_{\uparrow}/\Gamma_{\downarrow}$, and [62]
@@
$\gamma_{1}=\frac{2\pi}{\tau_{1}}=\Gamma_{\downarrow}+\Gamma_{\uparrow}.$ (23)
@@ A.2 Transmon Qubit Coupled to Several Baths
$\frac{p_{e}}{p_{g}}=\frac{\Sigma_{i}\Gamma_{\uparrow}^{(i)}}{\Sigma_{i}\Gamma_{\downarrow}^{(i)}}=\frac{\Sigma_{i}\gamma_{1}^{(i)}n\left(T^{(i)},\omega\right)}{\Sigma_{i}\gamma_{1}^{(i)}\left[n\left(T^{(i)},\omega\right)+1\right]}$
@@
$\gamma_{1}=\sum_{i}\left(\Gamma_{\downarrow}^{(i)}+\Gamma_{\uparrow}^{(i)}\right)=\sum_{i}\gamma_{1}^{(i)}\left[2\,n\left(\omega,T^{(i)}\right)+1\right].$ (24)
@@
$\bar{n}=\frac{\sum_{i}\Gamma_{\uparrow}^{(i)}}{\sum_{i}\left(\Gamma_{\downarrow}^{(i)}-\Gamma_{\uparrow}^{(i)}\right)}=\frac{\sum_{i}\gamma_{i}n\left(T^{(i)},\omega\right)}{\sum_{i}\gamma_{i}}=n\left(T_{\mathrm{eff}},\omega\right).$ (25)
@@
All considerations above can be generalized to the case of an arbitrary number of baths.
Ratio Calculation method Result Equation
1 2 3
@@ C.1 Mutual Dependencies
In order to find the relation between the errors of $A,B$ and $C$ and the measured temperature $T$ one can use logarithmic derivatives of the functions $A(T),B(T)$ and $C(T)$. In order to simplify the final expressions, in the low temperature limit expressions we can write
@@
where $x=\hbar\omega_{ge}/k_{B}T$. At low temperatures $x\rightarrow\infty$, Eqs. (40) and (41) give approximately equal errors, meaning that the relative errors of the temperatures $T_{B}$ and $T_{C}$ are similar in the relevant temperature range. That agrees very well with the data presented in this paper (see Sec. V and VI) and in Ref. [29].
@@ Appendix D Sample Fabrication
The next step was forming of Josephson junctions (JJs) (see Fig. 1 c). For that the sample was spincoated by a MMA/PMMA e-beam resist bilayer. A 800 nm thick layer of MMA EL-11 was applied by two steps of spincoating at 4000 rpm and baking at 160C for 2 min. Then a 200 nm layer of PMMA A4 was applied by spinning at 4000 rpm and baking at 160C for 10 min. Finally, an EBL step was performed to create a suspended mask for further JJ shadow deposition. The main and undercut doses for exposure of submicron structures was 950 $\mu$C/cm2 and 150 $\mu$C/cm2, respectively, with 1 nA beam current and 4 nm stepsize. The development of the mask was done in a two-step process: (i) development in MIBK:IPA 1:3 mixture for 25 s and (ii) in methylglycol:methanol 2:1 mixture for 15 s with a following rinsing in IPA for 30 s. After that the samples were loaded into an e-beam evaporator Plassys MEB700S2-III UHV for JJ deposition. First, the sample was pumped down to $5\cdot 10^{-7}$ mbar in the loadlock and an Ar milling was done to remove the resist residues at Ar pressure of $4\cdot 10^{-4}$ mbar with ion-beam gun (IBG) parameters of 60 mA beam current and 200 V beam voltage at angles ±$\pm 12^{\circ}$ for 30 s at each angle. Then the substrate was moved into the oxidation chamber and pumped to $1\cdot 10^{-7}$ mbar. After that the chambers were connected and Ti gettering was performed at a rate of 0.5 Å/s for 5 min. Then the first layer of 25 nm thick Al was deposited at angle $-12^{\circ}$ at 1 Å/s rate at pressure below $5\cdot 10^{-8}$ mbar in the oxidation chamber, after which the evaporation chamber was closed and an oxidation was done at a pressure of 18 mbar during 7 min. The second layer of 45 nm thick Al was deposited at $+12^{\circ}$ at 1 Å/s rate at a pressure below $1\cdot 10^{-7}$ mbar in the oxidation chamber. Finally, another oxidation was performed at pressure 20 mbar during 10 min. The lift-off procedure was done in hot acetone at $52^{\circ}$ for 3 h, then the sample was rinsed with IPA and dried with a N2 gun.
@@
Now the third step of EBL was done to create patches (see Fig. 1 c) for a galvanic connection between JJs and the Nb ground plane. The MMA/PMMA e-beam resist bilayer was applied as described above and after that the sample was loaded into the EBL tool. The patch exposure was done with the dose of 950 $\mu$C/cm2 with 30 nA beam current and 20 nm stepsize. The resist was developed by putting into MIBK:IPA 1:3 mixture for 30 s and then into methylglycol:methanol 2:1 for 30 s followed by rinsing in IPA for 30 s. For the patch deposition we again used the Plassys tool. Analogously to the previous step, the sample in the loadlock was pumped to $5\cdot 10^{-7}$ mbar and a stronger Ar milling at zero angle was done for 2 min to remove the surface Nb and Al oxides. The Ar pressure was $4\cdot 10^{-4}$ mbar and the IBG parameters were the following: 120 mA beam current, 400 V beam voltage. Then the sample was moved to the oxidation chamber and pumped to $1\cdot 10^{-7}$ mbar, after which a deposition of 150 nm of Al was performed at the rate of 2 Å/s and pressure below $1\cdot 10^{-7}$ mbar in the oxidation chamber with consecutive oxidation at 20 mbar during 10 min. Another lift-off step was done similarly to the one described above.

### v3

@@ V.3 Operation Speed
The presented method refers to the equilibrium type of quantum thermometry, meaning that the population measurement is performed after the qubit thermalization to the bath, which happens on time scales not shorter than the relaxation time $\tau_{1}$. In the presented experiment the time delay between the consecutive measurements, which is necessary for the qubit relaxation after the readout and thermalization of the qubit, was $\sim 10\,\tau_{1}$, which takes up most of the measurement time, and defines the speed of the response of the system. Pulse application and the qubit state readout can be limited to sub-microsecond range. Finally, the total measurement time is defined by the averaging number $t_{\mathrm{tot}}\sim N_{\mathrm{av}}\cdot 10\,\tau_{1}\simeq 10^{6}\tau_{1}$, with $N_{\mathrm{av}}\simeq 10^{5}$, meaning a time scale of seconds or minutes per point, depending on $\tau_{1}$.
V.4 Perspective Applications - On-Chip Thermometry of Mesoscopic Structures
Currently we do not consider qubit thermometers to be advantageous over well-established techniques such as calibrated thermometry of macroscopic objects. However, using them in applications similar to mesoscopic thermometers [79, 80, 81, 82, 12, 14], such as NIS-junctions, Coulomb-blockade thermometers (CBT), shot noise thermometers (SNT) or thermometers based on quantum dots may be perspective, especially, taking into account qubit’s minimal heat capacity (single photon needed for excitation) and low power readout (in single or few photons regime), which makes this thermometry highly noninvasive for the measured object.
In case of thermometry of a specific heat bath, coupled to a qubit, we can apply the same considerations as in Section IV.1. Let us assume, that the the effective environment is the object of our measurements. Then if the coupling in known, its temperature can be found from the saturation temperature of the qubit (see Eq. (21)). The latter is valid even in the case of the weak coupling and low temperatures of the object (Eq. (23)). If the coupling to the object is unknown, but the bath temperature is variable, the dependence $T_{\mathrm{eff}}\left(T_{\mathrm{env}}\right)$ can still be measured. However, in reality the situation can be complicated by presence of other parasitic heat baths, which could reduce the sensitivity of the qubit to the temperature of the object.
An example of possible realization of qubit thermometry of another mesoscopic object is coupling a normal metal resistor with well defined properties such as resistance $R$ and temperature $T$ capacitively or inductively to a qubit [83, 84, 85, 86, 87, 88]. In this case, as shown in Eq. (9), the qubit will be thermalized to the resistor if its relaxation rate to the resistor $\gamma_{R}$ will be much larger, than the total relaxation through all other channels, i.e. $\gamma_{R}\approx\gamma_{1}^{0}$, but low enough to allow for pulse measurements of the qubit, i.e. $1/\gamma_{R}\gtrsim 10\,\mu$s.
VI Conclusion
In this work we experimentally realized equilibrium quantum thermometry of the cryostat temperature, based on the population distribution measurements of the three lowest energy levels of a transmon qubit. The presented technique allowed for temperature measurements in the range of $60-200\,$mK, in which the qubit population follows the Maxwell-Boltzmann distribution. Experimental data can be well explained by a two-bath model, which accounts for presence of an effective environment with a fixed finite temperature, different from that of the cryostat. We provided analysis of the effective temperature of the qubit, coupled to both cryostat and the environment having different temperatures and relative coupling strength.
@@ VI Conclusion
We analyzed the working range of the qubit thermometer based on experimental data and numerical models and discussed possible ways to improve it. The temperature range of the thermometer was defined in the low temperature limit by the saturation temperature of $65-85\,$mK and above $200$ mK by quasiparticle suppression of the qubit relaxation time. The experimental data were compared to a numerical model, describing time evolution of the qubit population, which showed good agreement. We briefly discuss a prospective application of the qubit thermometer for measurements of on-chip mesoscopic heat baths. Finally, we discussed the accuracy of the employed technique and compared it with other experiments and theoretical limits. We believe, that this experimental platform allows for exploitation of quantum thermometry algorithms including nonequilibrium techniques for faster measurements and dynamical control of the thermometers for enhanced precision.
@@ Acknowledgements.
The authors are grateful to Gershon Kurizki, Mohammad Mehboudi and Luis Alberto Correa Marichal for fruitful discussions and Yu-Cheng Chang for assistance with measurements and fabrication. We acknowledge the provision of facilities and technical support by Aalto University at OtaNano - Micronova Nanofabrication Centre and OtaNano - Low Temperature Laboratory.
@@ Appendix A Relaxation and Excitation Rates for Coupling to a Thermal Bath
where $n=1/\left[\exp\left(\hbar\omega_{0}\beta\right)-1\right]$ is the Bose-Einstein distribution. The total decay rate in this case would be
@@ A.1 Effective Temperature of a Two-Level System
Let us consider a qubit with the ground state |$\ket{g}$ and excited state |$\ket{e}$, thermalized to a heat bath with temperature $T$. Then the population ratio of the qubit levels should follow the detailed balance principle $p_{e}/p_{g}=e^{-\hbar\omega\beta}=\Gamma_{\uparrow}/\Gamma_{\downarrow}$, and [90]
@@
$\gamma_{1}=\frac{1}{\tau_{1}}=\Gamma_{\downarrow}+\Gamma_{\uparrow}.$ (39)
@@ A.2 Transmon Qubit Coupled to Several Baths
$\frac{p_{e}}{p_{g}}=\frac{\Sigma_{i}\Gamma_{\uparrow}^{(i)}}{\Sigma_{i}\Gamma_{\downarrow}^{(i)}}=\frac{\Sigma_{i}\gamma^{(i)}n\left(T^{(i)},\omega\right)}{\Sigma_{i}\gamma^{(i)}\left[n\left(T^{(i)},\omega\right)+1\right]}$ (40)
@@
$\gamma_{1}=\sum_{i}\left(\Gamma_{\downarrow}^{(i)}+\Gamma_{\uparrow}^{(i)}\right)=\sum_{i}\gamma^{(i)}\left[2\,n\left(\omega,T^{(i)}\right)+1\right].$ (41)
@@
$\bar{n}=\frac{\sum_{i}\Gamma_{\uparrow}^{(i)}}{\sum_{i}\left(\Gamma_{\downarrow}^{(i)}-\Gamma_{\uparrow}^{(i)}\right)}=\frac{\sum_{i}\gamma^{(i)}n\left(T^{(i)},\omega\right)}{\sum_{i}\gamma^{(i)}}=n\left(T_{\mathrm{eff}},\omega\right).$ (42)
@@ C.1 Mutual Dependencies
In order to find the relation between the errors of $A,B$ and $C$ and the measured temperature $T$ one can use logarithmic derivatives of the functions $A(T),B(T)$ and $C(T)$. To simplify the final expressions, in the low temperature limit we can write
@@
where $x=\hbar\omega_{ge}/k_{B}T$. At low temperatures $x\rightarrow\infty$, Eqs. (57) and (58) give approximately equal errors, meaning that the relative errors of the temperatures $T_{B}$ and $T_{C}$ are similar in the relevant temperature range. That agrees very well with the data presented in this paper (see Section IV and V) and in Ref. [20].
@@ Appendix D Sample Fabrication
The next step was forming of Josephson junctions (JJs) (see Fig. 1 f). For that the samples was spincoated by a MMA/PMMA e-beam resist bilayer. A 800 nm thick layer of MMA EL-11 was applied by two steps of spincoating at 4000 rpm and baking at 160C for 2 min. Then a 200 nm layer of PMMA A4 was applied by spinning at 4000 rpm and baking at 160C for 10 min. Finally, an EBL step was performed to create a suspended mask for further JJ shadow deposition. The main and undercut doses for exposure of submicron structures was 950 $\mu$C/cm2 and 150 $\mu$C/cm2, respectively, with 1 nA beam current and 4 nm stepsize. The development of the mask was done in a two-step process: (i) development in MIBK:IPA 1:3 mixture for 25 s and (ii) in methylglycol:methanol 2:1 mixture for 15 s with a following rinsing in IPA for 30 s. After that the samples were loaded into an e-beam evaporator Plassys MEB700S2-III UHV for JJ deposition. First, the samples were pumped down to $5\cdot 10^{-7}$ mbar in the loadlock and an Ar milling was done to remove the resist residues at Ar pressure of $4\cdot 10^{-4}$ mbar with ion-beam gun (IBG) parameters of 60 mA beam current and 200 V beam voltage at angles ±$\pm 12^{\circ}$ for 30 s at each angle. Then the substrate was moved into the oxidation chamber and pumped to $1\cdot 10^{-7}$ mbar. After that the chambers were connected and Ti gettering was performed at a rate of 0.5 Å/s for 5 min. Then the first layer of 25 nm thick Al was deposited at angle $-12^{\circ}$ at 1 Å/s rate at pressure below $5\cdot 10^{-8}$ mbar in the oxidation chamber, after which the evaporation chamber was closed and an oxidation was done at a pressure of 18 mbar during 7 min. The second layer of 45 nm thick Al was deposited at $+12^{\circ}$ at 1 Å/s rate at a pressure below $1\cdot 10^{-7}$ mbar in the oxidation chamber. Finally, another oxidation was performed at pressure 20 mbar during 10 min. The lift-off procedure was done in hot acetone at $52^{\circ}$ for 3 h, then the samples were rinsed with IPA and dried with a N2 gun.
@@
Now the third step of EBL was done to create patches (see Fig. 1 f) for a galvanic connection between JJs and the Nb ground plane. The MMA/PMMA e-beam resist bilayer was applied as described above and after that the samples were loaded into the EBL tool. The patch exposure was done with the dose of 950 $\mu$C/cm2 with 30 nA beam current and 20 nm stepsize. The resist was developed by putting into MIBK:IPA 1:3 mixture for 30 s and then into methylglycol:methanol 2:1 for 30 s followed by rinsing in IPA for 30 s. For the patch deposition we again used the Plassys tool. Analogously to the previous step, the samples in the loadlock were pumped to $5\cdot 10^{-7}$ mbar and a stronger Ar milling at zero angle was done for 2 min to remove the surface Nb and Al oxides. The Ar pressure was $4\cdot 10^{-4}$ mbar and the IBG parameters were the following: 120 mA beam current, 400 V beam voltage. Then the samples were moved to the oxidation chamber and pumped to $1\cdot 10^{-7}$ mbar, after which a deposition of 150 nm of Al was performed at the rate of 2 Å/s and pressure below $1\cdot 10^{-7}$ mbar in the oxidation chamber with consecutive oxidation at 20 mbar during 10 min. Another lift-off step was done similarly to the one described above.